• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name brads

    Log entry time 15:20:55 on February 12, 2009

    Entry number 260672

    keyword=BB Cerenkov ADC weirdness

    I noticed that the pedestals for PMTs 2, 10, 18 were really lousy. Basically a smear across all ADC channels with a big overflow pileup.

    The signals look "OK" on the scope. I powered off the HV to PMT 2 and took a run. Pedestal went back to normal (ie. 1-2 bin wide spike). Powered it back on -- pedestal got a little wider (as expected with the PMT powered up), but now looked quite good.

    "Power cycled" the HV for PMT channel 10 and that channel was fixed.

    "Power cycled" HV for PMT 18 and problem remained. Unplugged all inputs from slot 17 ADC and problem remained. Power cycled the VME crate, problem went away -- all pedestals look normal again. Ick.

    There is already a v792Reset() call in the Download section of the CRL, so we're doing everything we can do in software at the start of each run. I guess I'll add some pedestal histos to the online gui to monitor the situation.

    A copy of this log entry has been emailed to: brads